4. Suppose you titrate 50.00 mL of 1.20 M pyridine, C5H5N, with 0.634 M HCl.

(a) How many moles of pyridine are present in the original pyridine solution?

(b) What is the initial pH?

(c) What volume (in mL) of HCl solution is required to reach the equivalence point?

(d) What is the pH at the equivalence point?

(e) What is the pH of the solution after the addition of 104 mL of HCl?
Answer #1

pH formulas for salts weak acid + strong baise salt tho pH at { pk w tp ka +Log e] strong acid + weak base salt tho pH = [one litre of Molarity = no- of moles of solufe present in the solution is called Molarity -number of moles & of solute MolariÇHEN: Kg = value = 1+2x109 paba-log kb = -10g (17X109) = 8-77 moles of estan X1000 nolarity of - Stan volume Imp saltan 1000크622ac 2 . - 이키 나2세에)를u)(-2-04×169) 2 C) | -- xot O-000 ioss 3 아s | -- xo + - Coo lossa 22 Xe CoHS = o. oooo LS2761 M PoFe ACSH-N molarity = 12 M volume soml=0-054 Mat no of moles = 12X0-05 20.06. at equivalence point moles of moles of onof mole ç HNFor this type of salt pt formula pH = t[ pk w pkb_logo pkwaly pkb = 8.77 c=0.414823 M pHaţ[14-8-77-log (9.414823)] pH= 2.8060SHSN: molarity = R2 M volumea soml=0.052 M= 2 no of moles=kz X0-05 20-06 | Hel: Molarity = 0.634M volume = 104m) 20.1042 no o